OpenStudy (anonymous):

find all angles that solve the expression sec(t)=4

OpenStudy (anonymous):

@satellite73 Is H(x) A.O.S X=-2

OpenStudy (anonymous):

i think you need a calculator for this

OpenStudy (anonymous):

@D1Prospect yes

OpenStudy (anonymous):

Ok Thanks Man

OpenStudy (anonymous):

\[\sec(t)=4\\ \cos(t)=\frac{1}{4}\\ t=\arccos(\frac{1}{4})\]

OpenStudy (anonymous):

that is only one, not all, but there is no way for you to know this number without a calculator